OpenStudy (anonymous):

symbol of the cation. i know CuO=Cu^2+ but what is Ag2O

OpenStudy (anonymous):

well if O is ^-2 then Ag must have what charge to balance?

OpenStudy (callisto):

since CuO is neutral, charge of O = 0-2 =-2 Consider Ag2O, it is also neutral, hence charge of Ag = [0-(-2)] /2 = +1. Therefore, it is Ag^+

OpenStudy (anonymous):

Charge of Ag = +1